OpenStudy (anonymous):

An acute angle θ is in a right triangle with cos θ =(5)/(6) . What is the value of sec θ? and An acute angle θ is in a right triangle with sin θ = (5)/(6) . What is the value of cot θ?

OpenStudy (ajprincess):

\[\sec\theta=\frac{ 1 }{ \cos\theta }\] can u find the value of \(sec\theta\) @TGNG?

OpenStudy (anonymous):

When doing the first problem you must remember that cos and sec are opposite. Cos = adj/hyp and sec=hyp/adj. Therefore, if cos is 5/6, then sec is 6/5. It works the same way with the second part.

OpenStudy (anonymous):